OpenStudy (anonymous):

Probability!!!!!!!!!!!!!!!!!!!! A bag has 2 yellow, 7 red, and 1 green marble. What is the probability of picking a red marble, replacing it, and then picking another red marble?

OpenStudy (anonymous):

7/10 chances

OpenStudy (foreverandalways13):

weelll it would be out of ten marbles and you have 7 red and trying to get 2 red out of ten so meaning you have 7/10 chance to get red

OpenStudy (amistre64):

thats P(red) we want P(red) and P(red)

OpenStudy (foreverandalways13):

hope thats not confusing...

OpenStudy (amistre64):

... and its not 7/10

OpenStudy (amistre64):

the prob of getting a red on the first try is: 7/10 replacing it and trying to pull out a red again is: 7/10 P(red) and P(red) = 7/10 * 7/10

OpenStudy (anonymous):

you have the same probability both times you reach in - 7/10 7/10 x 7/10 = 49/100